# Demystifying the Compression of Mixture-of-Experts Through a Unified Framework

### v2

@@ Abstract
Scaling large language models has revolutionized the performance across diverse domains, yet the continual growth in model size poses significant challenges for real-world deployment. The Mixture of Experts (MoE) approach addresses this by dynamically selecting and activating only a subset of experts, significantly reducing computational costs while maintaining high performance. However, MoE introduces potential redundancy (e.g., parameters) and extra costs (e.g., communication overhead). Despite numerous compression techniques developed for mitigating the redundancy in dense models, the compression of MoE remains under-explored. We first bridge this gap with a cutting-edge unified framework that not only seamlessly integrates mainstream compression methods but also helps systematically understand MoE compression. This framework approaches compression from two perspectives: Expert Slimming which compresses individual experts and Expert Trimming which removes structured modules. Within this framework, we explore the optimization space unexplored by existing methods, and further introduce aggressive Expert Trimming techniques, i.e., Layer Drop and Block Drop, to eliminate redundancy at larger scales. Based on these insights, we present a comprehensive recipe to guide practitioners in compressing MoE effectively. Extensive experimental results demonstrate the effectiveness of the compression methods under our framework and the proposed recipe, achieving a 6.05×6.05$\times$6.05 × speedup and only $20.0$GB memory usage while maintaining over $92\%$ of performance on Mixtral-8$\times$7B. Code is released at https://github.com/DaizeDong/Unified-MoE-Compression.
@@ 1 Introduction
While scaling large language models has shown exceptional performance across various domains [1, 2, 3], the increasing model size poses significant challenges in real-world deployments [4, 5] due to excessive computational demands and associated costs. The Mixture of Experts (MoE) [6], which selectively activates a subset of parameters during inference, offers a promising solution to reduce these computational burdens. Additionally, integrating MoE with Large Language Models (LLMs) [7, 8] has been shown to enhance performance further.
Despite these advancements, existing MoE models still exhibit redundancies that increase deployment costs. Common MoE approaches often replicate feedforward layers multiple times, leading to parameter redundancy. For example, He et al. [9] observed that expert parameters could be compressed through parameter-sharing, Liu et al. [10] revealed extremely high representational similarity between experts, while Lu et al. [11] noted that not all experts are essential, with some being safely removable. Additionally, MoE models introduce communication overheads compared to dense models [12, 13, 14] in distributed settings, as tokens must be distributed among different experts and their outputs subsequently combined.
In this paper, we investigate the issues through the lens of compression to enhance the efficiency of MoE [15, 16]. Specifically, we propose a unified framework that integrates two complementary perspectives: (1) Expert Slimming that compresses individual experts and (2) Expert Trimming that structurally removes experts. In the case of Expert Slimming, network pruning [17, 18] and quantization [19, 20] are the most widely used techniques, where quantization is hardware-friendly for implementations. In terms of existing Expert Trimming, we utilize Expert Drop [11, 21], which removes unimportant experts.
@@
This framework facilitates us to systematically understand the efficiency issue of MoE and identify the new design space to further improve the performance. With the framework, we first reveal the efficiency bottlenecks of existing methods, specifically the communication between experts and costly computation. Based on this insight, we propose aggressive Expert Trimming methods to enhance MoE efficiency. Specifically, to mitigate communication and computation costs, we present Layer Drop that removes all experts in an MoE layer, i.e., the entire MoE layer. Additionally, given the computation-intensive nature of the attention mechanism within transformer blocks, we propose Block Drop, which removes both attention layers and MoE layers. We use similarity-based metrics to demonstrate the feasibility of Layer Drop and Block Drop. Surprisingly, these two coarse-grained methods outperform fine-grained Expert Drop by a large margin in balancing performance and efficiency. Furthermore, since Expert Slimming can be seamlessly integrated with Expert Trimming, we further utilize the unified framework to propose a comprehensive compression recipe that integrates both strategies.
@@
Experimental results on two widely-used MoE models (i.e., Mixtral-8$\times$7B [7] and DeepSeek-MoE-16B [8]) demonstrate the effectiveness of our methods. For Expert Slimming, quantization performs best, since it significantly compresses by a large margin practically while maintaining comparable performance. For Expert Trimming, Expert Drop significantly reduces the memory usage but only accelerates the inference marginally; Layer Drop and Block Drop result in both faster inference speed and less memory usage while maintaining comparable performance. The combined strategy of Expert Slimming and Expert Trimming achieves a 6.05$\times$ speedup and only 20.0GB memory usage while maintaining over 92% of the performance on Mixtral-8$\times$7B. These findings offer valuable insights for enhancing the efficiency of MoE models.
@@ 2 Related Work
Mixture of Experts. The Mixture of Experts (MoE) is a kind of dynamic neural network with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation [22, 23, 24]. The potential of sparse activation in MoE is subsequently exploited by [6] for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision [25] and language [26, 27, 28] scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (LLMs) [7, 8, 29, 30], achieving superior scaling laws at low costs [31]. Further investigations emerge in developing improved expert structures [32, 33, 8], router designs [34, 35, 36, 37], and training strategies [38, 39, 10, 40], propelling the continuous evolution on the representation capability and computational efficiency of MoE models. Despite the success, MoE also suffers from efficiency issues. For instance, MoE replicates the experts, significantly increasing the parameter budget [9]. On the other hand, adopting multiple experts to process input tokens introduces communication costs and enhances latency [13, 41].
Compression Methods. The escalating size of large language models presents considerable hurdles for their practical implementation. Consequently, a range of efficient methods has emerged to address the high-cost deployment issues. Among them, model quantization [5, 42, 43, 44] and network pruning [4, 45, 46, 47] are widely utilized. Model quantization reduces the precision of neural network weights to lower bits [19], while network pruning [17] removes redundant parameters or architectures. Although these methods have shown promising results on dense models, they lack consideration for the inductive bias inherent in MoE. To bridge this gap, Expert Drop, as proposed in studies like [21, 11], addresses the unique nature of MoE by removing unimportant experts. By eliminating redundant experts, the MoE architecture becomes more compact and can be deployed at a lower cost. However, while Expert Drop leads to a more compact architecture, it may also lead to non-negligible performance drop and rely on post-training procedures for recovery.
3 Preliminary
@@ 3.1 Mixture of Experts
A standard Mixture of Experts layer comprises a set of $n$ experts $\left\{\boldsymbol{E}_{1},\boldsymbol{E}_{2},\dots,\boldsymbol{E}_{n}\right\}$ associated with corresponding weights $\{\boldsymbol{W}_{1},\boldsymbol{W}_{2},\dots,\boldsymbol{W}_{n}\}$ and a router $\boldsymbol{G}$. The router determines which experts to activate for a given input $\boldsymbol{x}$ by calculating selection scores $\boldsymbol{G}(\boldsymbol{x})\in\mathbb{R}^{n}$ for all experts. It then selects the top $k$ experts with the highest scores, resulting in a sparse activation pattern. The input $\boldsymbol{x}$ is then sent to the selected experts, and their outputs are combined into a weighted sum based on the selection scores provided by the router. This process can be described by the following equations:
$\mathcal{K}=\mathrm{TopK}(\mathrm{Softmax}(\boldsymbol{G}(\boldsymbol{x})),k),$ (1)
$\boldsymbol{y}=\sum\nolimits_{i\in\mathcal{K}}{\boldsymbol{G}(\boldsymbol{x})_{i}\cdot\boldsymbol{E}_{i}(\boldsymbol{x}|\boldsymbol{W}_{i})},$ (2)
where $\mathcal{K}$ denotes the indices of selected experts, $\boldsymbol{G}(\boldsymbol{x})_{i}$ represents the selection score for the $i$-th expert, and $\boldsymbol{E}_{i}(\boldsymbol{x})$ is the output from the $i$-th expert. In transformer models, the MoE layer is often used as a replacement for the feed-forward network (FFN). In this context, each expert functions as an independent FFN module, enhancing the model’s capacity without a proportional increase in computational cost [48].
Challenges. While MoE models have demonstrated strong performance across various tasks [7, 8], they also encounter significant deployment issues: On one hand, MoE replicates expert networks multiple times, increasing model size and memory costs. On the other hand, the complexity of managing multiple expert networks and the associated communication between them leads to increased latency and slower inference speeds.
@@ 3.2 Overview of Previous Compression Methods
To address the challenges arising from MoE, we investigate several mainstream and state-of-the-art efficient methods for MoE.
Pruning: For a MoE layer that includes multiple experts $\{\boldsymbol{E}_{i}\}_{i=1}^{n}$ with corresponding weights $\{\boldsymbol{W}_{i}\}_{i=1}^{n}$, pruning uses binary masks $\{\boldsymbol{M}_{i}\}_{i=1}^{n}$ to selectively disable parts of weights:
$\hat{\boldsymbol{W}}_{i}=\boldsymbol{M}_{i}\odot\boldsymbol{W}_{i}.$ (3)
Masking can exhibit unstructured [49, 50], semi-structured, or structured properties, each resulting in different sparsity patterns. Unstructured sparsity often exhibits superior performance, while semi-structured sparsity offers a balance between performance and efficiency. Structured sparsity is hardware-friendly, but its performance is currently deemed insufficient.
Quantization: Unlike pruning, which involves masking out unimportant parameters, quantization transforms weights into low-bit representations, thereby reducing memory requirements.
$\hat{\boldsymbol{W}}_{i}=\rm{Quant}(\boldsymbol{W}_{i}),$ (4)
where “$\mathrm{Quant}$” denotes the quantization function. Quantization reduces the complexity of data types but does not impact the FLOPs (floating point operations) or the total number of parameters.
Expert Drop: Different from fine-grained pruning and quantization, Expert Drop entails the removal of expert networks, based on the observation that not all experts are equally important [11, 21]. Given expert-wise importance scores $\boldsymbol{S}$ (e.g., the routing scores, $\boldsymbol{S}(\boldsymbol{E}_{i})=\boldsymbol{G}(\boldsymbol{x})_{i}$), Expert Drop retains only the experts with the highest $n^{\prime}$ scores:
$\mathcal{T}^{\prime}=\mathrm{TopK}(\boldsymbol{S}(\{\boldsymbol{E}_{i}\}_{i=1}^{n}),n^{\prime}),$ (5)
$\boldsymbol{E}\leftarrow\{\boldsymbol{E}_{i}\}_{i\in\mathcal{T}^{\prime}},\quad\boldsymbol{G}\leftarrow\{\boldsymbol{G}_{i}\}_{i\in\mathcal{T}^{\prime}}.$ (6)
@@
Here, $\mathcal{T}^{\prime}$ denotes the subset of the original expert indices $\mathcal{T}=\{1,2,\dots,n\}$. Expert Drop  reduces FLOPs conditionally: when $\mathcal{T}^{\prime}$ contains more than or equal to $k$ indices, MoE still utilizes the top $k$ experts for each input; otherwise, it uses all remaining experts. While this approach reduces communication between experts, the resulting speedup is not always significant.
Others: Other efficient methods include Low-Rank compression [51, 52], which decomposes model weights into smaller weights. In this paper, we mainly focus on the widely-used methods (i.e., pruning, quantization, dropping), deferring discussion on other methods for future exploration.
4 A Unified View on MoE Compression
In this section, we propose a general framework that unifies the efficient methods on MoE. This framework enables us to systematically understand the efficiency issues of MoE models and identify new design opportunities to further improve performance.
4.1 Unified View for MoE Compression
Existing compression methods address these challenges by focusing on either the internal compression of individual experts or the structural simplification of the model. Inspired by this insight, we unify the MoE compression into two complementary perspectives: Expert Slimming that compresses individual experts (e.g., Pruning and Quantization), and Expert Trimming that directly removes experts (e.g., Expert Drop). In doing so, we can better understand their contributions and limitations. An overview of these perspectives is illustrated in Figure 1.
The compression of individual experts (Expert Slimming) focuses on the transformation and reduction of expert weights, denoted as $\boldsymbol{W}$. We utilize a transformation function $f(\boldsymbol{W})$ to represent this process. The transformation function $f(\boldsymbol{W})$ can be understood as a general mapping that applies various compression techniques to the weights of the model. For example, in pruning, $f(\boldsymbol{W})$ could be a function that sets a subset of the weights to zero. In quantization, $f(\boldsymbol{W})$ might reduce the precision of the weights from 32-bit floats to 8-bit integers. On the other hand, Expert Trimming  deals with compressing structured modules by selecting and retaining only a subset of the experts, denoted as $\mathcal{T}^{\prime}$. This is represented by the transformation $\mathcal{T}\leftarrow\mathcal{T}^{\prime}$. Methods like Expert Drop, which selectively drops entire experts or layers, are examples of this approach. By integrating these two perspectives, we can derive a general form for efficient MoE models. The compression within and across experts can be expressed as follows:
$\boldsymbol{y}=\sum\nolimits_{i\in\mathcal{T}^{\prime}}\boldsymbol{G}_{i}\cdot\boldsymbol{E}_{i}(\boldsymbol{x}|f(\boldsymbol{W}_{i})).$ (7)
Table 1: Summary of Compression Methods. “✓” means effective and “✗” means ineffective, while “❍” represents conditionally effective, depending on specific settings and environments.
Hierarchy Method Formulation Parameter Memory FLOPs Speedup
Pruning $\boldsymbol{M}\odot\boldsymbol{W}$
Expert Slimming Quantization Quant($\boldsymbol{W}$)
Expert $\mathcal{T}\leftarrow\mathcal{T}^{\prime}$

### v3

# Towards Efficient Mixture of Experts: A Holistic Study of Compression Techniques

@@ Abstract
Scaling large language models has driven remarkable advancements across various domains, yet the continual increase in model size presents significant challenges for real-world deployment. The Mixture of Experts (MoE) architecture offers a promising solution by dynamically selecting and activating only a subset of experts during inference, thus substantially reducing computational costs while preserving high performance. Despite these benefits, MoE introduces new inefficiencies, such as excessive parameters and communication overhead. In this work, we present a holistic study of compression techniques for Mixture of Experts to enhance both efficiency and scalability. While recent efforts have focused on Expert Trimming, which reduces the number of experts, these approaches still suffer from considerable communication and computational costs. To address this, we propose more aggressive strategies, such as Layer Drop, which removes entire MoE layers, and Block Drop, which eliminates transformer blocks. Surprisingly, these aggressive pruning techniques not only preserve model performance but also substantially improve computation and memory efficiency. Furthermore, beyond Expert Trimming, we also introduce Expert Slimming, which compresses individual experts to further boost performance and can be seamlessly integrated with Expert Trimming. Extensive experimental results demonstrate the effectiveness of our proposed methods — Layer Drop and Block Drop — along with the comprehensive recipe that integrates Expert Slimming and Expert Trimming, achieving a 6.05×6.05$\times$6.05 × speedup with $77.1$% reduced memory usage while maintaining over $92\%$ of performance on Mixtral-8$\times$7B. Our code is released at https://github.com/CASE-Lab-UMD/Unified-MoE-Compression.
@@ 1 Introduction
While scaling large language models has shown exceptional performance across various domains (Ramesh et al., 2021; OpenAI, 2024; Team, 2024a), the increasing model size poses significant challenges in real-world deployments (Sun et al., 2023; Frantar et al., 2022) due to excessive computational demands and associated costs. The Mixture of Experts (MoE) (Shazeer et al., 2017), which selectively activates a subset of parameters during inference, offers a promising solution to reduce these computational burdens. Additionally, integrating MoE with Large Language Models (LLMs) has been shown to further enhance performance (Jiang et al., 2024; Dai et al., 2024).
Despite these advancements, MoE models still suffer from significant redundancies that increase deployment costs. Standard MoE implementations replicate feed-forward layers across multiple experts, resulting in heavily parameterized models. For instance, Mixtral-8$\times$7B (Jiang et al., 2024) contains 47B parameters, but only 13B parameters are activated per token, leading to substantial GPU memory consumption and limited scalability. In addition, replicating experts often introduces redundant experts. For example, He et al. (2023) observed that expert parameters could be compressed through parameter sharing. Similarly, Lu et al. (2024) noted that not all experts are essential, suggesting that some can be safely removed. These findings underscore the potential for compressing MoE models to improve efficiency without sacrificing effectiveness.
@@
In this paper, we first investigate the Expert Trimming based compression techniques, which reduce the number of experts to enhance the efficiency of MoE (Cheng et al., 2020; Liang et al., 2021). The most prevalent approach for Expert Trimming is Expert Drop, which scores each expert and drops the less important ones (Lu et al., 2024; Muzio et al., 2024). While Expert Drop reduces model size, it does not eliminate costly computations within the MoE layer and complex communication among experts, leading to negligible improvements in the inference speed. To this end, we propose more aggressive Expert Trimming methods to enhance MoE efficiency. Specifically, to mitigate communication and computation costs, we present Layer Drop that removes the entire MoE layer. Additionally, given the computation-intensive nature of the attention mechanism within transformer blocks, we further propose Block Drop, which removes the whole transformer blocks. We use similarity-based metrics to demonstrate the feasibility of Layer Drop and Block Drop. Surprisingly, these two coarse-grained methods outperform fine-grained Expert Drop by a large margin in balancing performance and efficiency. Additionally, with small-scale post-finetuning, the compressed models can be further optimized to achieve near-original performance.
Beyond removing experts, we also explore Expert Slimming, which focuses on compressing individual experts. Techniques such as network pruning (Han et al., 2016; Zhu & Gupta, 2017) and quantization (Jacob et al., 2017; Nagel et al., 2021) have proven effective for the model compression, with quantization being particularly suitable for hardware acceleration. By integrating Expert Slimming with Expert Trimming, we propose a unified framework for compressing MoE models that further maximizes efficiency gains while maintaining strong performance.
@@
Our experimental results on representative MoE models, Mixtral-8$\times$7B (Jiang et al., 2024) and DeepSeek-MoE-16B (Dai et al., 2024), demonstrate the effectiveness of our proposed methods. For Expert Trimming, Expert Drop significantly reduces the memory usage but it provides only marginal improvements in inference speed. In contrast, Layer Drop and Block Drop significantly accelerate inference and reduce memory usage while maintaining comparable performance to the original models. The combined strategy of Expert Trimming and Expert Slimming results in a 6.05$\times$ speedup with only 22.8% memory usage (20.0GB) while maintaining over 92% of the original performance on Mixtral-8$\times$7B. The findings offer valuable insights for enhancing the efficiency of MoE models. Additionally, post-finetuning allows compressed models to recover most of their original performance, resulting in a minimal 0.6% performance gap compared to the uncompressed DeepSeek-MoE-16B model.
In summary, by conducting a holistic study on compressing Mixture of Experts, our key contributions are:
  • We extend Expert Trimming to a higher architectural with Layer Drop and Block Drop, significantly enhancing computation and memory efficiency while preserving model performance.
  • We integrate Expert Trimming with Expert Slimming to further achieve efficiency gains without compromising performance.
  • Extensive experimental results demonstrate the effectiveness of our proposed methods, achieving a 6.05×6.05$\times$6.05 × speedup and reducing memory usage to just $20.0$ GB, all while maintaining over $92\%$ of performance on Mixtral-8$\times$7B.
@@ 2 Related Work
Mixture of Experts
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation (Jacobs et al., 1991; Jordan & Jacobs, 1994). The potential of sparse activation in MoE is subsequently exploited by Shazeer et al. (2017) for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision (Riquelme et al., 2021) and language (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022) scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (LLMs) (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024a; Zhu et al., 2024; Team, 2024b), achieving superior scaling laws at low computational costs (Clark et al., 2022). Further investigations emerge in developing improved expert structures (Gururangan et al., 2022; Rajbhandari et al., 2022; Dai et al., 2024), router designs (Lewis et al., 2021; Roller et al., 2021; Zhou et al., 2022), and training strategies (Shen et al., 2023; Chen et al., 2022), propelling the continuous evolution on the representation capability and computational efficiency of MoE models. Despite the success, MoE also suffers from efficiency issues. For instance, MoE replicates the experts, significantly increasing the parameter budget (He et al., 2023). On the other hand, adopting multiple experts to process input tokens introduces communication costs and enhances latency (Song et al., 2023; Xue et al., 2024b).
Compression Methods
The increasing size of large language models presents considerable challenges for their practical implementation. Consequently, a range of efficient methods has emerged to address the implementation issues. Among them, model quantization (Frantar et al., 2022; Lin et al., 2024) and network pruning (Sun et al., 2023; Frantar & Alistarh, 2023) are widely utilized. Model quantization reduces the precision of neural network weights to lower bits (Jacob et al., 2017), while network pruning (Han et al., 2016) removes redundant parameters or architectures. Although these methods have shown promising results on dense models, they lack consideration for the inductive bias inherent in MoE. To bridge this gap, Expert Drop, as proposed in studies like (Muzio et al., 2024; Lu et al., 2024), addresses the unique nature of MoE by removing unimportant experts. By eliminating redundant experts, the MoE architecture becomes more compact and can be deployed at a lower cost. However, while Expert Drop leads to a more compact architecture, it may also lead to non-negligible performance drop and rely on post-training for recovery.
3 Preliminaries
@@ 3.1 Mixture of Experts
A Mixture of Experts (MoE) layer consists of a collection of $n$ experts, $\{\bm{E}_{1},\bm{E}_{2},\dots,\bm{E}_{n}\}$, each associated with weights $\{\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{n}\}$, and a router $\bm{G}$ that dynamically selects the most relevant experts for a given input $\bm{x}$. The router computes selection scores, $\bm{G}(\bm{x})\in\mathbb{R}^{n}$, for all experts and selects the top $k$ experts, resulting in a sparse activation pattern. The input $\bm{x}$ is processed by the selected experts, and their outputs are combined into a weighted sum based on the router’s scores. This process is mathematically expressed as:
$\mathcal{K}=\mathrm{TopK}(\mathrm{Softmax}(\bm{G}(\bm{x})),k),$ (1)
$\bm{y}=\sum\nolimits_{i\in\mathcal{K}}{\bm{G}(\bm{x})_{i}\cdot\bm{E}_{i}(\bm{x}|\bm{W}_{i})},$ (2)
where $\mathcal{K}$ denotes the indices of selected experts, $\bm{G}(\bm{x})_{i}$ represents the selection score for the $i$-th expert, and $\bm{E}_{i}(\bm{x})$ is the output from the $i$-th expert. In transformer models, the MoE layer usually replaces the feed-forward network (FFN). In this context, each expert functions as an independent FFN module, enhancing the model’s capacity without a proportional increase in the computational cost (Vaswani et al., 2017).
Challenges
While MoE models have demonstrated strong performance across various tasks (Jiang et al., 2024; Dai et al., 2024), they also encounter significant deployment challenges. On one hand, MoE models replicate multiple expert networks, inflating the model size and memory usage. For instance, Mixtral-8$\times$7B has a total of 47B parameters and requires 87.7GB of memory for deployment, though only 13B parameters are activated for each token. On the other hand, the communication required to manage multiple expert networks increases latency and slows down inference speed, especially in distributed environments (Song et al., 2023; Yu et al., 2024).
@@ 3.2 Overview of Previous Compression Methods
To address the efficiency challenges, we first review several mainstream and state-of-the-art compression techniques for MoE models.
Pruning: Pruning reduces the number of active parameters by selectively disabling parts of the model’s weights. In an MoE layer with $n$ experts $\{\bm{E}_{i}\}_{i=1}^{n}$ and corresponding weights $\{\bm{W}_{i}\}_{i=1}^{n}$, pruning introduces binary masks $\{\bm{M}_{i}\}_{i=1}^{n}$ to deactivate certain weights:
$\hat{\bm{W}}_{i}=\bm{M}_{i}\odot\bm{W}_{i}.$ (3)
Pruning can be unstructured (Lee et al., 2021; Bai et al., 2022), semi-structured, or structured. Unstructured sparsity tends to yield the best performance, semi-structured sparsity strikes a balance between efficiency and performance, and structured sparsity, while hardware-friendly, often results in lower performance.
Quantization: Unlike pruning, which involves masking out unimportant parameters, quantization reduces memory usage by converting model weights to lower-bit representations. For MoE layers, quantization is applied in the following way:
$\hat{\bm{W}}_{i}=\text{Quant}(\bm{W}_{i}),$ (4)
where “$\mathrm{Quant}$” denotes the quantization function. Quantization decreases the computation and memory consumption without reducing FLOPs or the total number of parameters, making it particularly advantageous for hardware acceleration.
Expert Drop: Different from fine-grained pruning and quantization, Expert Drop entails the removal of expert networks, based on the observation that not all experts are equally important (Lu et al., 2024; Muzio et al., 2024). Given expert-wise importance scores $\bm{S}$ (e.g., the routing scores, $\bm{S}(\bm{E}_{i})=\bm{G}(\bm{x})_{i}$), Expert Drop retains only the experts with the highest $n^{\prime}$ scores:
$\mathcal{T}^{\prime}=\mathrm{TopK}(\bm{S}(\{\bm{E}_{i}\}_{i=1}^{n}),n^{\prime}),$ (5)
$\bm{E}\leftarrow\{\bm{E}_{i}\}_{i\in\mathcal{T}^{\prime}},\quad\bm{G}\leftarrow\bm{G}_{i\in\mathcal{T}^{\prime}}.$ (6)
Here, $\mathcal{T}^{\prime}$ denotes the subset of the original expert indices $\mathcal{T}=\{1,2,\dots,n\}$. Expert Drop  reduces FLOPs conditionally: when $\mathcal{T}^{\prime}$ contains more than or equal to $k$ indices, MoE still utilizes the top $k$ experts for each input; otherwise, it uses all remaining experts. While this approach reduces communication between experts, the resulting speedup is usually insignificant when maintaining acceptable performance.
Other Compression Techniques: Other methods, such as low-rank decomposition (Li et al., 2024b; a), aim to compress model weights into smaller matrices, further reducing memory and computational costs. In this work, we primarily focus on the widely-used methods (pruning, quantization, and Expert Drop), leaving a more detailed exploration of these additional methods for future research.
4 A Holistic Study of MoE Compression Techiniques
@@
In this section, we propose a general framework that unifies various compression methods for MoE. This framework provides a comprehensive understanding of MoE model efficiency issues and helps identify new design spaces for further performance improvements.
4.1 Overview
Existing MoE compression methods primarily address two types of inefficiencies: structural redundancies in the overall architecture and internal redundancies within individual experts. To address both issues, we categorize these methods into two complementary perspectives: Expert Trimming that focuses on removing structured components (e.g., experts, layers, and blocks), and Expert Slimming that compresses individual experts through techniques like pruning or quantization. An overview of these compression perspectives is illustrated in Figure 1.
Expert Trimming  deals with compressing structured modules by selecting and retaining only a subset of the experts, denoted as $\mathcal{T}\textquoteright$ ’. This is represented by the transformation $\mathcal{T}\leftarrow\mathcal{T}\textquoteright$ ’. Methods like Expert Drop, which selectively drops unimportant experts, are examples of this approach. On the other hand, the compression of individual experts (Expert Slimming) focuses on the transformation and reduction of expert weights, denoted as $\bm{W}$. We utilize a transformation function $f(\bm{W})$ to represent this process. The transformation function $f(\bm{W})$ can be understood as a general mapping that applies various compression techniques to the weights of the model. For example, in pruning, $f(\bm{W})$ could be a function that sets a subset of the weights to zeros. In quantization, $f(\bm{W})$ might reduce the precision of the weights from 32-bit floats to 8-bit integers. By integrating these two perspectives, we can derive a general form for efficient MoE models. The compression within and across experts can be expressed as follows:
$\bm{y}=\sum\nolimits_{i\in\mathcal{T}^{\prime}}\bm{G}_{i}\cdot\bm{E}_{i}(\bm{x}|f(\bm{W}_{i})).$ (7)
In the following sections, we will elaborate on Expert Trimming and Expert Slimming, respectively.
Table 1: Summary of Compression Methods. “✓” means effective, indicating that the method performs well as intended. “✗” means ineffective, where the method fails to optimize the specified metric. “❍” represents conditionally effective, depending on specific settings and environments.
Method Formulation Parameter Memory FLOPs Speedup
Expert Drop $\mathcal{T}\leftarrow\mathcal{T}^{\prime}$